Solve the fraction below using number line

1. 4 2/3 / 1 1/6


2. 1 1/4 / 2/3


3. 2/5 / 1/3


4. 3/4 / 5/6



1
Expert's answer
2021-08-03T13:55:01-0400

Solution (1)

"4\frac{2}{3}, 1\frac{1}{6}"


or "\frac{14}{3},\frac{7}{6}"





Solution (2)

"1\frac{1}{4},\frac{2}{3}"


or "\frac{5}{4},\frac{2}{3}"





Solution (3)

"\frac{2}{5},\frac{1}{3}"





Solution (4)

"\frac{3}{4},\frac{5}{6}"